four times the greater of two consecutive integers is 1 more than five times the lesser number. what are the integers?
Answer by mananth(16946)

let x & x+1 be the integers
..
4 times the greater = 4(x+1)
1 mor than 5 times the sammler 5x+1
4(x+1)=5x+1
4x+4 =5x+1
5x-4x=3
x= 3
3 & 4 are the integers
...
m.ananth@hotmail.ca